# LLM $\TITLELLM for Large-Scale Optimization Model Auto-Formulation$: A Lightweight Few-Shot Learning Approach

### v1

@@
LLM $\TITLELLM for Large-Scale Optimization Model Auto-Formulation$: A Lightweight Few-Shot Learning Approach
@@ 2 Department of Industrial Systems Engineering and Management, National University of Singapore
3 School of Information Management and Engineering, Shanghai University of Finance and Economics
@@ 7 SC Johnson College of Business, Cornell University
Large-scale optimization is a key backbone of modern business decision-making. However, building these models is often labor-intensive and time-consuming. We address this by proposing LEAN-LLM-OPT, a LightwEight AgeNtic workflow construction framework for LLM-assisted large-scale OPTimization auto-formulation. LEAN-LLM-OPT takes as input a problem description together with associated datasets and orchestrates a team of LLM agents to produce an optimization formulation. Specifically, upon receiving a query, two upstream LLM agents dynamically construct a workflow that specifies, step-by-step, how optimization models for similar problems can be formulated. A downstream LLM agent then follows this workflow to generate the final output. Leveraging LLMs’ text-processing capabilities and common modeling practices, the workflow decomposes the modeling task into a sequence of structured sub-tasks and offloads mechanical data-handling operations to auxiliary tools. This design alleviates the downstream agent’s burden related to planning and data handling, allowing it to focus on the most challenging components that cannot be readily standardized. Extensive simulations show that LEAN-LLM-OPT, instantiated with GPT-4.1 and the open source gpt-oss-20B, achieves strong performance on large-scale optimization modeling tasks and is competitive with state-of-the-art approaches. In addition, in a Singapore Airlines choice-based revenue management use case, LEAN-LLM-OPT demonstrates practical value by achieving leading performance across a range of scenarios. Along the way, we introduce Large-Scale-OR and Air-NRM, the first comprehensive benchmarks for large-scale optimization auto-formulation. The code and data of this work is available at https://github.com/CoraLiang01/lean-llm-opt.
@@ 1 Introduction
As modern businesses grow in scale and complexity, organizations increasingly rely on advanced data analytics and optimization tools to enhance decision-making and operational efficiency. From supply chain networks to financial planning and resource allocation, many business sectors face the challenge of extracting information from data and formulating intricate optimization problems on top of it (Chen and Graves 2021, Kang et al. 2022, Dang et al. 2021, Deng et al. 2023). However, as observed in many cases, in order to apply these techniques, companies and organizations usually rely on highly skilled experts to understand the problems, process data, formulate the optimization models, and solve them. Currently, the above works typically require significant manual effort and can be costly, making it a major bottleneck in fully leveraging the power of optimization. Therefore, researchers start to explore automated modeling of optimization problems in an end-to-end manner, i.e., from the natural language problem descriptions and data to optimization formulations and the corresponding programming codes (Ramamonjison et al. 2022, 2023).
@@
Recently, advancements in Large Language Models (LLMs) open a new avenue for automated optimization modeling (Simchi-Levi et al. 2025). Built on LLMs’ remarkable ability in understanding and processing unstructured information and natural languages, several studies have been focusing on utilizing LLMs to formulate and/or solve optimization problems based on text descriptions. This facilitates the application of optimization techniques by automating expert-level modeling, allowing more non-experts to handle tricky optimization problems through interfacing with LLMs. As initial attempts, several works examine how to prompt pretrained LLMs to formulate and/or solve optimization problems directly (Xiao et al. 2023, AhmadiTeshnizi et al. 2024, Li et al. 2023, Chen et al. 2024, Wasserkrug et al. 2025). Perhaps not surprisingly, using prompting techniques only without any modification to the LLMs may lead to unsatisfactory accuracy. Hence, researchers turn to fine-tuning-based approaches, which modify the parameters of open-source LLMs based on task-specific datasets, to improve their performance (Ma et al. 2024, Huang et al. 2024, 2025, Yang et al. 2024, Jiang et al. 2025, Astorga et al. 2024).
@@
While fine-tuning-based approaches significantly improve LLMs’ modeling accuracy, prior works primarily focus on small-scale inputs (i.e., the problem) and outputs (i.e., the optimization model). Meanwhile, a practical optimization model typically involves a large number of variables, and in order to correctly define its objective and constraints, one usually needs to process data stored in large-scale datasets (e.g., CSV files). However, as we can see in Table 5 of Section 4, when turning to large-scale problems [^1] For inputs, we refer to problem instances whose data has to be stored separately in datasets as large-scale. For outputs, we define the scale based on the number of variables needed (small-size: <20<20 variables, medium-size: 209920-99 variables, large-size: $\geq 100$ variables). In this work, we focus mainly on large-scale inputs but allow the scale of the output to vary., the modeling accuracy of state-of-the-art fine-tuned models (e.g., ORLM proposed in Huang et al. (2025)) and closed-source models (e.g., GPT-5.2) can quickly deteriorate. This thus raises a critical question, can we design LLM-based methods for efficient large-scale optimization modeling from natural language inputs? For this, several major challenges persist:
@@
  • Challenge 1. Long Input Processing and Reasoning: In large-scale optimization modeling, LLMs must navigate long inputs that combine diverse problem descriptions with associated datasets, making standardization and automation particularly challenging. Although recent LLMs are designed to handle lengthy and complex reasoning tasks, prior work has reported that reasoning performance can degrade substantially as input length increases (Levy et al. 2024, Li et al. 2025, Du et al. 2025). To examine whether this phenomenon also arises in the optimization modeling setting, Figure 1 reports the performance of several state-of-the-art models on our curated dataset Large-Scale-OR (introduced in Section 2). We observe that the accuracies of most models decrease noticeably as the input size grows. In particular, the modeling accuracies of Gemini 3 Pro and GPT-5.2, the flagship models by Google and OpenAI, both drop below 50% when the number of input tokens exceeds 800—well below their advertised input limits (on the order of 100,000 to 1,000,000 tokens);
@@
  • Challenge 2. Inadequate Training Resource: A tempting solution for the above is to continue with fine-tuning approaches. While these can work for small-scale problems, generalizing the training pipelines to large-scale optimization modeling can be unrealistic. This is because large-scale optimization problems can be much more versatile. For example, by altering a small portion of the problem description or the datasets (e.g., split a dataset into multiples, change the names and order of the rows or columns), one can get a totally different problem instance and a new labeled may also be needed. Consequently, one may need to construct a prohibitively large set of inputs and labeled outputs to apply fine-tuning approaches.
@@ 1.1 Main Contributions and Key Messages
To address the above challenges, we propose LEAN-LLM-OPT, a novel LightwEight AgeNtic workflow construction framework for LLM-assisted large-scale OPTimization model auto-formulation. Leveraging human expertise in optimization modeling, LEAN-LLM-OPT takes a problem description and associated datasets as input and coordinates three LLM agents to output the optimization formulation (together with the Python programming code to solve the model). Our main contributions can be summarized as follows:
@@
    Innovative Agentic Workflow Construction Approach: For every input, LEAN-LLM-OPT first uses a classification agent to determine the type of the problem (e.g., network revenue management, resource allocation). Then, as the key innovation of LEAN-LLM-OPT (see Figure 3 for a graphical illustration), a workflow generation agent dynamically constructs a workflow that illustrates how to use customized tools and reasoning to formulate optimization models for this type of problem. The workflow specifies which components of the datasets require focused attention and guides the agent to gradually form better understanding of the target model. Finally, a model generation agent follows this workflow to understand the problem, extract relevant data with our customized tools, and generate the final answer. Through these steps, LEAN-LLM-OPT successfully decomposes a long and complicated reasoning task into a process with multiple structured steps, enabling an LLM to execute one after another with its reasoning power (we provide a more in-depth discussion on its advantages in Section 3.3);
@@
    High-Quality Dataset Curation: As an intermediate step, we turn data acquired from Kaggle (Kaggle 2025) into a set of problem instances. Each problem instance contains a user query (problem description and associated datasets) and label (the problem type and optimization model). From these, we create a reference dataset Ref-Data of 96 entries for the agents to use as reference. Additionally, we build a 101-entry testing dataset Large-Scale-OR, serving as a benchmark to reflect diverse practical business applications. Different from existing benchmark datasets that focus mostly on small-size problem instances (less than 20 variables), Large-Scale-OR puts emphasis on medium-size (20 to 99 variables, constitutes 26% of the dataset) and large-size (at least 100 variables, constitutes 50% of the dataset) problem instances. Both Ref-Data and Large-Scale-OR span a variety of applications across different domains. To ensure our method is generalizable, we purposely make the instances in Ref-Data substantially different from those in Large-Scale-OR in terms of length and semantics (see Section 2.3);
@@
    Extensive Numerical Simulations: We conduct numerical simulations across different datasets to evaluate the performance of LEAN-LLM-OPT built on GPT-4.1 and gpt-oss-20B (an open source model from OpenAI that can address cost and privacy concerns). In Large-Scale-OR, both of our model achieve an overall accuracy of over 80%, significantly outperforming other models such as ORLM, Gemini 3 Pro (Google’s latest reasoning model), and GPT-5.2 (OpenAI’s latest reasoning model). We also test the performance of our models in other widely-used small-scale benchmarks such as NL4OPT (Ramamonjison et al. 2023), IndustryOR (Huang et al. 2025), Mamo (Easy and Complex) (Huang et al. 2024). The results show that our models continue to achieve high overall accuracy;
@@
    Singapore Airlines Use Case and Ablation Study: Finally, we apply our techniques to a Singapore Airlines’ use case that concerns its fare type capacity allocation and flight network planning. We focus on the sales-based linear programming (SBLP) formulation (Gallego et al. 2015) using Singapore Airlines’ simulated datasets. We start from the fare type capacity allocation problem and show that our approach can achieve a leading accuracy of over 86% in Air-NRM-CA, a benchmark that consists of 15 problem instances. Next, we consider the network planning problem (i.e., joint fare type capacity allocation and flight scheduling). Our results show that LEAN-LLM-OPT can consistently achieve an optimality gap less than 15% in Air-NRM-NP, a benchmark that consists of 21 problem instances. These outperform Gemini 3 Pro and GPT-5.2 in almost all instances. Through our ablation study, we can clearly see that both the agentic workflow and the data-handling tools are of great importance for the high modeling accuracy. These findings not only underscore the effectiveness but also showcase the practicality of our design.
@@ 2.1 Reference Dataset Construction
Each instance in Ref-Data can be expressed using a tuple $(q,t,g,f,m)$. The user’s query $q$ includes the problem descriptions and the corresponding datasets, the problem type is denoted as $t$, and the optimization model $m$ is generated by our experts. Besides, we also use $g$ and $f$ to additionally capture the related data category (i.e., the type of the data that is relevant to the optimization model) and the details of the relevant data. We note that some of the instances in Ref-Data can be of small-scale as they are helpful in problem classification. An example of a problem instance in Ref-Data is given below [^2] To test the generalization capability of our method and avoid overfitting, we purposely keep the instances in Ref-Data simple. We shall report the comparison between Ref-Data and Large-Scale-OR different in length and semantics in Section 2.3..
@@ 2.3 Textual Comparisons between Ref-Data and Large-Scale-OR
Besides, we also compute the average token counts of Ref-Data and Large-Scale-OR using the o200k_base tokenizer from the tiktoken library and the results is presented in Table 3. As we can see, the average token count of the instances in Ref-Data is much smaller compared to the average token count of the instances in Large-Scale-OR.
Table 3: Average token counts of Ref-Data and Large-Scale-OR
Ref-Data Large-Scale-OR
422 47269
Together, we can conclude that the instances in Ref-Data are substantially different from those in the benchmarks. Also, the instances in Large-Scale-OR are significantly more challenging than those in Ref-Data. Hence, in order to perform well across all the benchmarks, a method has to be able to generalize well across both the semantic and input scale dimensions.
@@ 3 LEAN-LLM-OPT Framework
We are now ready to present the design details of LEAN-LLM-OPT. LEAN-LLM-OPT builds model for large-scale optimization models through the collaboration of three LLM agents as illustrated in Figure 3. Specifically, after receiving an input, which includes a problem description and the corresponding datasets, the classification agent identifies the problem type and outputs it to the workflow generation agent. Then, the workflow generation agent retrieves the most relevant example of the same type from Ref-Data, and builds the workflow that demonstrates how to build the model for the example. Subsequently, this workflow is provided to the model generation agent. The latter adheres to the well-defined, step-by-step workflow to generate the optimization formulation. In what follows, we use the below large-scale query as a running example to introduce the three agents.
@@ 3.1 Problem Classification
After receiving the query, the classification agent tries to identify the type of the input problem (e.g., one of those in Table 1). This is to make sure the subsequent workflow and agents can capture the unique characteristics of the problem to formulate the model.
@@
  • Question: In this step, a sample problem description (different from the user’s one) whose type will be determined following the remaining steps is provided.
@@
After receiving the input, the classification agent follows the demo to output the problem type by executing each of the steps. For our running example, the type of the input is identified as “Resource Allocation Problem”. As we can see, every step above require certain level of reasoning, but each of them is kept to minimal so that it becomes manageable for the agent.
Remark 1 (The Benefit of Classification and Advantages of Reasoning).
We note that the classification agent is not strictly required as LEAN-LLM-OPT can perform well even without it (see the numerical experiment results in Table 5). However, using a classification agent can help customize the subsequent workflow for commonly encountered problem types, and hence, freeing up reasoning bandwidth for the downstream agents to further improve performance.
It is also worth pointing out that one may consider directly using RAG without LLM reasoning. However, as we show in Section D of the appendix, the classification accuracy can be negatively impacted if so. This underscores the importance of leveraging the reasoning capability of the LLMs.

### v2

# LLM $\TITLELLM for Large-Scale Optimization Model Auto-Formulation$: Bridging Flexibility and Standardization via Agentic Workflow

@@
LLM $\TITLELLM for Large-Scale Optimization Model Auto-Formulation$: Bridging Flexibility and Standardization via Agentic Workflow
@@ 2 Department of Industrial Systems Engineering and Management, National University of Singapore
3 Department of Analytics & Operations, Imperial Business School
@@ 7 SC Johnson College of Business, Cornell University
Large-scale optimization is a key backbone of modern business decision-making. However, building these models is often labor-intensive and time-consuming. We address this by proposing LEAN-LLM-OPT, a LightwEight AgeNtic workflow construction framework for LLM-assisted large-scale OPTimization model auto-formulation. LEAN-LLM-OPT takes as input a problem description together with associated datasets and orchestrates a team of LLM agents to produce an optimization formulation. Specifically, upon receiving a query, two upstream LLM agents dynamically construct a workflow that specifies, step-by-step, how optimization models for similar problems can be formulated; A downstream LLM agent then follows this workflow to generate the final output. The agentic workflow leverages common modeling practices to standardize the modeling process into a sequence of structured sub-tasks, offloading mechanical data-handling operations to auxiliary tools. This reduces the LLM’s burden in planning and data handling, allowing us to exploit its flexibility to address unstructured components. Extensive simulations show that LEAN-LLM-OPT, instantiated with GPT-4.1 and the open source gpt-oss-20B, achieves strong performance on large-scale optimization modeling tasks and is competitive with state-of-the-art approaches. In addition, in a Singapore Airlines choice-based revenue management use case, LEAN-LLM-OPT demonstrates practical value by achieving leading performance across a range of scenarios. Along the way, we introduce Large-Scale-OR and Air-NRM, the first comprehensive benchmarks for large-scale optimization auto-formulation. The code and data of this work is available at https://github.com/CoraLiang01/lean-llm-opt.
@@ 1 Introduction
As modern businesses grow in scale and complexity, organizations increasingly rely on advanced optimization techniques to enhance decision-making and operational efficiency. From supply chain networks to financial planning and resource allocation, many business sectors face the challenge of extracting information from data and formulating intricate optimization problems on top of it (Chen and Graves 2021, Kang et al. 2022, Dang et al. 2021, Deng et al. 2023). As observed in many cases, in order to apply these techniques, companies and organizations usually rely on highly skilled experts to understand the problems, process data, and formulate the optimization models. Currently, the above works typically require significant manual effort and can be costly, making it a major bottleneck in fully leveraging the power of these techniques. Therefore, researchers start to explore automated modeling of optimization problems in an end-to-end manner, i.e., from the natural language problem descriptions and data to optimization formulations and the corresponding programming codes (Ramamonjison et al. 2022, 2023).
@@
Recently, advancements in Large Language Models (LLMs) open a new avenue for automated optimization modeling (Simchi-Levi et al. 2025). Built on LLMs’ remarkable ability in understanding and processing unstructured information and natural languages, several studies have been focusing on utilizing LLMs to formulate and/or solve optimization problems based on text descriptions. This facilitates the application of optimization techniques by automating expert-level modeling, allowing more non-experts to handle tricky optimization problems through interfacing with LLMs. As initial attempts, several works examine how to prompt pretrained LLMs to formulate and/or solve optimization problems directly (Xiao et al. 2023, AhmadiTeshnizi et al. 2024, Li et al. 2023, Chen et al. 2024, Wasserkrug et al. 2025). Perhaps not surprisingly, using prompting techniques only without any modification to the LLMs may lead to unsatisfactory accuracy. Hence, researchers turn to fine-tuning approaches, which modify the parameters of open source LLMs based on labeled datasets, to improve their performance (Ma et al. 2024, Huang et al. 2024, 2025, Yang et al. 2024, Jiang et al. 2025, Astorga et al. 2024).
@@
While fine-tuning approaches significantly improve LLMs’ modeling accuracy, prior works primarily focus on small-scale inputs and outputs, where the data can be directly embedded into the problem description and the output only has a few variables. Meanwhile, a practical optimization model typically involves a large number of variables, and in order to correctly define its objective and constraints, the input usually has a large volume of data that has to be stored separately in datasets. However, when turning to large-scale problems [^1] We refer to problem instances whose data has to be stored separately in datasets as large-scale. For the corresponding model, we define the scale based on the number of variables needed (small-size: <20<20 variables, medium-size: 209920-99 variables, large-size: $\geq 100$ variables). In this work, we focus mainly on large-scale inputs but allow the scale of the output to vary. See Section 2.3 for a summary statistics for the scale of the inputs of our benchmark., the modeling accuracies of state-of-the-art fine-tuned models and closed-source commercial models can quickly deteriorate (see, e.g., Table 6 of Section 4). This thus raises a critical question, can we design LLM-based methods for efficient large-scale optimization modeling from natural language inputs? For this, several major challenges persist:
@@
  • Challenge 1. Long Input Processing and Reasoning: In large-scale optimization modeling, LLMs must navigate long inputs that combine diverse problem descriptions with associated datasets, making standardization and automation particularly challenging. Although recent LLMs are designed to handle lengthy and complex reasoning tasks, prior work has reported that reasoning performance can already degrade substantially simply because input length increases (Levy et al. 2024, Li et al. 2025, Du et al. 2025). To examine whether this phenomenon also arises in the optimization modeling setting, Figure 1 reports the performance of several state-of-the-art models on our curated dataset Large-Scale-OR (introduced in Section 2). We observe that the accuracies of most models decrease noticeably as the input size grows. In particular, the modeling accuracies of Gemini 3 Pro and GPT-5.2, the flagship models by Google and OpenAI, both drop below 50% when the number of input tokens exceeds 800–well below their advertised input limits (on the order of 100,000 to 1,000,000 tokens);
@@
  • Challenge 2. Inadequate Training Resource: A tempting solution for the above is to continue with fine-tuning approaches so that they become more flexible and adept for a wider range of problems. While these can work for small-scale problems, generalizing the training pipelines to large-scale optimization modeling can be unrealistic. This is because large-scale optimization problems can be much more versatile. For example, by altering a small portion of the problem description or the datasets (e.g., split a dataset into multiples, add or delete some rows and/or columns in the datasets), one may get a totally different problem instance and a new labeled may also be needed. Consequently, one may need to construct a prohibitively large set of inputs and labeled outputs to apply fine-tuning approaches.
@@ 1.1 Main Contributions and Key Messages
To address the above challenges, we propose LEAN-LLM-OPT, a novel LightwEight AgeNtic workflow construction framework for LLM-assisted large-scale OPTimization model auto-formulation. LEAN-LLM-OPT takes a problem description and associated datasets as input and coordinates three LLM agents to output the optimization formulation (together with the Python programming code to solve the model). Our main contributions can be summarized as follows:
@@
    Innovative Agentic Workflow Construction Approach: For every input, LEAN-LLM-OPT first uses a classification agent to determine the type of the problem (e.g., network revenue management, resource allocation). Then, as the key innovation of LEAN-LLM-OPT (see Figure 3 for a graphical illustration), a workflow generation agent dynamically constructs a workflow with a reference problem to provide an illustrative procedure for applying customized tools and structured reasoning to formulate optimization models for this problem class. The workflow specifies which components of the datasets require focused attention and guides the agent to gradually form better understanding of the target model. Finally, a model generation agent follows this workflow to understand the problem, extract relevant data with our customized tools, and generate the final answer. Through these steps, LEAN-LLM-OPT successfully decomposes a long and complicated reasoning task into a process with multiple structured steps, enabling an LLM to execute one after another with its reasoning power (we provide a more in-depth discussion on its advantages in Section 3.3);
@@
    High-Quality Dataset Curation: As an intermediate step, we turn data acquired from Kaggle (Kaggle 2025) into a set of problem instances. Each problem instance contains a user query (problem description and associated datasets) and label (the problem type and optimization model). From these, we build a 101-entry testing dataset Large-Scale-OR, serving as a benchmark to reflect diverse practical business applications. At the same time, we also create a reference dataset Ref-Data, whose entries are substantially different from those in Large-Scale-OR in terms of semantics and scale (see Section 2.3 for a detailed measurement), for the agents to use as references. Different from existing benchmark datasets that focus mostly on small-size problem instances (less than 20 variables), Large-Scale-OR puts emphasis on medium-size (20 to 99 variables, constitutes 26% of the dataset) and large-size (at least 100 variables, constitutes 50% of the dataset) problem instances. Both Ref-Data and Large-Scale-OR span a variety of applications across different domains;
@@
    Extensive Numerical Simulations: We conduct numerical simulations across different datasets to evaluate the performance of LEAN-LLM-OPT built on GPT-4.1 and gpt-oss-20B (an open source model from OpenAI). In Large-Scale-OR, both of our models achieve an overall accuracy of above 76%, substantially outperforming other models such as ORLM, Gemini 3 Pro, and GPT-5.2. We also test the performance of our models in other widely-used small-scale benchmarks such as NL4OPT (Ramamonjison et al. 2023), IndustryOR (Huang et al. 2025), Mamo (Easy and Complex) (Huang et al. 2024). The results show that our models continue to achieve high modeling accuracies;
@@
    Singapore Airlines Use Case and Ablation Study: Finally, we apply our techniques to a Singapore Airlines’ use case that concerns its fare type capacity allocation and flight network planning. We focus on the sales-based linear programming (SBLP) formulation (Gallego et al. 2015) using Singapore Airlines’ simulated datasets. We start from the fare type capacity allocation problem and show that our approach can achieve leading accuracices in Air-NRM-CA, a benchmark that consists of 15 problem instances. Next, we consider the network planning problem (i.e., joint fare type capacity allocation and flight scheduling). Our results show that LEAN-LLM-OPT can consistently achieve a small optimality gap in Air-NRM-NP, a benchmark that consists of 21 problem instances. These outperform the competing methods in almost all instances. Through our ablation study, we can clearly see that both the agentic workflow and the data-handling tools are of great importance for the high modeling accuracy. These findings not only underscore the effectiveness but also showcase the practicality of our design.
@@ 2.1 Reference Dataset Construction
Each instance in Ref-Data can be expressed using a tuple $(q,t,g,f,m)$. The user’s query $q$ includes the problem descriptions and the corresponding datasets, the problem type is denoted as $t$, and the optimization model $m$ is generated by our experts. Besides, we also use $g$ and $f$ to additionally capture the related data category (i.e., the type of the data that is relevant to the optimization model) and the details of the relevant data. We note that some of the instances in Ref-Data can be of small-scale as they are helpful in problem classification. An example of a problem instance in Ref-Data is given below [^2] To test the generalization capability of our method and avoid overfitting, we purposely keep the instances in Ref-Data simple. We shall report the comparison between Ref-Data and Large-Scale-OR in terms of length and semantics in Section 2.3..
@@ 2.3 Textual Comparisons between Ref-Data and Large-Scale-OR
Besides, we also compute the average token counts of Ref-Data, Large-Scale-OR, and other benchmarks using the o200k_base tokenizer from the tiktoken library and the results is presented in Table 3. As we can see, the average token count of the instances in Ref-Data is much smaller compared to the average token count of the instances in Large-Scale-OR. The scale of inputs in Large-Scale-OR is also much larger than those in other benchmarks.
Table 3: Average token counts of Ref-Data and different benchmarks
Ref-Data Large-Scale-OR NL4OPT MAMO Easy MAMO Complex IndustryOR
422 47269 121 217 457 266
Together, we can conclude that the instances in Ref-Data are substantially different from those in the benchmarks. Also, the instances in Large-Scale-OR are significantly larger in terms of token counts than those in Ref-Data. Hence, in order to perform well across all the benchmarks, a method has to be able to generalize well across both the semantic and input scale dimensions.
@@ 3 LEAN-LLM-OPT Framework
We are now ready to present the design details of LEAN-LLM-OPT. LEAN-LLM-OPT builds model for large-scale optimization problems through the collaboration of three LLM agents as illustrated in Figure 3. Specifically, after receiving an input, which includes a problem description and the corresponding datasets, the classification agent identifies the problem type and outputs it to the workflow generation agent. Then, the workflow generation agent retrieves the most relevant example of the same type from Ref-Data, and builds the workflow that demonstrates how to build the model for the example. Subsequently, this workflow is provided to the model generation agent. The latter adheres to the well-defined, step-by-step workflow to generate the optimization formulation. In what follows, we use the below large-scale query as a running example to introduce the three agents.
@@ 3.1 Problem Classification
After receiving the query, the classification agent tries to identify the type of the input problem (e.g., one of those in Table 1). This is to make sure the subsequent agents can capture the unique characteristics of the problem to generate the workflow and formulate the model.
@@
  • Question: In this step, a sample problem description (different from the user’s one) whose type will be determined following the remaining steps is provided;
@@
After receiving the input, the classification agent follows the demo to output the problem type by executing each of the steps. For our running example, the type of the input is identified as “Resource Allocation Problem” (although the groundtruth is “Mixture”). As we can see, every step above require certain level of reasoning, but each of them is kept to minimal so that it becomes manageable for the agent.
We point out that the incorporation of this classification agent is not strictly required, yet it can positively impact the modeling accuracy. We provide a discussion in the forthcoming Remark 2.